Find the condition on ‘a’ & ‘b’ so that the two tangents drawn to the parabola y^2 = 4ax from a point are normals to the parabola x^2 = 4by.


Grade:11

1 Answers

Himanshu Sharma
11 years ago

first find the slope of tangent to y^2=4ax

diff it to find he slope

2ydy=4adx

dy/dx=2a/y    (ii)

similarly find slope of tangent to x^2=4by

2xdx=4bdy

dy/dx=x/2b

slope of normal=-1/slope of tangent

hence slope of normal=-2b/x   -(i)

but (i)and (ii) are same

hence    -2b/x=2a/y

by+ax=0

 

by+ax=0